OpenStudy (anonymous):

Determine whether the problems should be solved using the formula for an arithmetic sequence, arithmetic series, geometric sequence, or geometric series. Explain your answer.do not need to solve.

OpenStudy (anonymous):

1. Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August? 2. A local grocery store stacks the soup cans in such a way that each row has 2 fewer cans than the row below it. If there are 32 cans on the bottom row, how many total cans are on the bottom 14 rows? 3. A major US city reports a 12% increase in decoration sales during the yearly holiday season. If decoration sales were 8 million in 1998, how much did the city report in total decoration sales by the end of 2004? 4. A fireplace contains 46 bricks along its bottom row. If each row above decreases by 4 bricks, how many bricks are on the 12th row?

OpenStudy (anonymous):

no one wants to help me :(

OpenStudy (anonymous):

Sara Idk how to answer this but ask either @dpalnc or @ParthKohli or @Hero.. either one of them might answer your question.. lol.. If they're BUSY or NOT.. just ask them..

OpenStudy (anonymous):

1. In February He deposited : 5 In March : 5*2 = 10 In April : 10*2 = 20 In May : 20*2 = 40 In June : 40*2 = 80 In July : 80*2 = 160 In August : 160*2 = 320

OpenStudy (anonymous):

i don't need to solve it

OpenStudy (anonymous):

It is a Geometric Sequence...

OpenStudy (anonymous):

Common Ratio = 10/5 = 20/10 = 40/20 = 2

OpenStudy (ash2326):

@sara1234please post one question at a time, tell me which one you don't get

OpenStudy (anonymous):

i just need to know what formulas i need to use to solve each of them don't need help solving

OpenStudy (ash2326):

cool, let's start with first one

OpenStudy (ash2326):

Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August? Feb=5 Next month it'll get doubled March=10 April=20 and so on we see the amount is in the sequence 5, 10, 20, 40.... so it's a geometric sequence with common ratio 2. Now you just need to figure out the amount in august feb= 1st month aug= 7th month Use the nth term formula \[a_n=a_1r^[n-1} here a_n is amount in august a_1= amount in feb r=2 tell me what's n here?

OpenStudy (anonymous):

The second one is Arithmetic Sequence with a common difference of two as each row on above is getting added by cans..

OpenStudy (ash2326):

@sara1234 do you get this?

OpenStudy (anonymous):

i don't need to solve it @ash2326

OpenStudy (ash2326):

But you did understand, didn't you?